# Heterophily-Aware Fair Recommendation using Graph Convolutional Networks

### v3

@@ 1 Introduction
Graph-based recommendation has become more popular in recent years due to the existence of a bipartite graph in the nature of recommender systems [5, 14, 23, 27, 28, 31]. In this bipartite graph, nodes are divided into two sets: users and items, with edges representing ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [43, 9, 11].
@@ 3 Preliminaries
In this section, we introduce preliminaries widely used in the paper, including task formulation, graph neural networks, the Light Graph Convolution model and the heterophily property of graphs.
@@ 3.2 Graph neural networks
Here, $D$ is the diagonal degree matrix that has the same dimension as the adjacency matrix $A$ of the graph, $D_{ii}=\sum_{j}A_{ij}$, $\tilde{D}=D+I$, $\tilde{A}=A+I$, $I\in R^{\left(|U|+|I|\right)\times\left(|U|+|I|\right)}$ is the identity matrix, $\sigma\left(\cdot\right)$ is an activation function, $H^{\left(k\right)}$ is the matrix of node embeddings (representations) in the layer $k$ and $M^{\left(k\right)}$ denotes a layer-specific trainable weight matrix. To compute the embedding of each node in the graph, a graph neural network constructs a graph, which is a rooted tree and is usually called node’s computational graph [32].
@@ 4.2 Fairness-aware attention
$h_{u}^{\left(k\right)}=\sum_{i\in N\left(u\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}\times\lambda\times\sigma(h_{u}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)})}h_{i}^{\left(k-1\right)},$ (8)
$h_{i}^{\left(k\right)}=\sum_{u\in N\left(i\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}\times\lambda\times\sigma(h_{i}^{\left(k-1\right)}\cdot h_{u}^{\left(k-1\right)})}h_{u}^{\left(k-1\right)},$ (9)
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\lambda$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. Therefore, the result of the sigmoid function for long-tail items will be close to zero, and for short-head items, it will be close to one. By incorporating the sigmoid of dot products along with the above-mentioned normalization term, long-tail items achieve a smaller denominator than short-head items. This ensures that larger embeddings, which are more similar to each other, are divided by larger values and shrink more, while smaller embeddings are divided by smaller values and thus shrink less. Therefore messages from long-tail items (such as $i$) receive a higher importance coefficient than those from short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients.
@@ 4.3 Heterophily feature weighting
$w_{ui}^{\left(k\right)}=\lambda\times\sigma\left(\left(h_{u}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)}\right)W^{\left(k\right)}\right),$ (10)
@@
4.4 Model training
$L_{BPR}=\sum_{\left(u,i,j\right)\in T}-\ln\sigma\left(\tilde{y}_{ui}-\tilde{y}_{uj}\right)+\beta||\Theta||^{2},$ (13)
4.5 The algorithm
@@
Algorithm 1 One batch training process.
1:Input: Graph $G\left(V,E\right)$, the number of layers $K$.
2:Output: Updated model parameters.
3:$X\leftarrow$ Xavier initialization $\triangleright$ Initialize input features
4:for $k=1$ to $K$ do
6:end for
7:$H\leftarrow X$triangleright Initialize embedding matrix
8:$L\leftarrow 0$triangleright Initialize BPR loss
9:$H\leftarrow$ FairEmbeddingGeneration($G,H,K$)
11:     Sample a negative item $j$ uniformly at random from $I\setminus O_{u}^{+}$
12:     $\tilde{y}_{ui}\leftarrow H\left[u\right]\cdot H\left[i\right]$
13:     $\tilde{y}_{uj}\leftarrow H\left[u\right]\cdot H\left[j\right]$
14:     $L\leftarrow L+\left(-\ln\sigma\left(\tilde{y}_{ui}-\tilde{y}_{uj}\right)+\beta||\Theta||^{2}\right)$
16:end for
17:return $\Theta$
@@
1:Input: Graph $G\left(V,E\right)$; input features matrix $X\in\mathbb{R}^{|V|\times d}$; number of layers $K$; weight matrices $W^{\left(k\right)}\in\mathbb{R}^{1\times d},\forall k\in\{1,\ldots,K\}$; parameter $\lambda$$\cdot$
@@
8:              $w_{vi}^{\left(k\right)}\leftarrow\lambda\times\sigma(\begin{bmatrix}s\end{bmatrix}\times W^{\left(k\right)})$triangleright By $\begin{bmatrix}s\end{bmatrix}$, we mean the $1\times 1$ matrix, consisting of scalar $s$.
@@ 5.1.1 Datasets
We examine the algorithms over six well-known datasets: Epinions [^1] https://snap.stanford.edu/data/soc-Epinions1.html [42], Amazon-Movies [^2] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html [13], Amazon-Electronics2 [13], Amazon-Health2 [13], Amazon-Beauty2 [13] and Amazon-CDs2 [13]. For the amazon review datasets, we adopt a $10$-core setting, filtering out users and items with less than $10$ interactions. In the case of Epinions, we utilize the data released by APDA  [42]. Table 1 depicts the statistics of the datasets.
@@
Dataset #users #items #interactions Density
Electronics 20242 11589 347393 0.148%
Epinions 11496 11656 327942 0.245%
Beauty 1340 733 28798 2.93%
Health 2184 1260 55076 2%
Movies 33326 21901 958986 0.131%
CDs 15592 16184 445412 0.176%
@@ 5.1.3 Evaluation metrics
Here, $SRC\left(\cdot,\cdot\right)$ calculates Spearman’s rank correlation, $\tilde{O}_{u}^{+}$ represents items in the test profile of user $u$, $rank_{u}\left(x\right)$ signifies the ranking of item $x$ that the model predicts for user $u$, $U_{i}$ is a set of users that item $i$ already exists in their test profiles and $d_{x}$ is the degree of node $x$. While PRU investigates items’ position bias for individual users’ recommendation lists, PRI considers item-side fairness across all users. A large value of the later may result in long-tail items having limited opportunities to become popular [44]. In general, the methods’ performance in promoting fairness improves as the values of these two criteria get closer to 00.
@@
Here, $REL_{N}$ is a sorted list of the $N$ most relevant items in ascending order, $r(i)$ is the relevance of the $i^{th}$ ranked item, which is either 00 or $1$ respectively for irrelevant and relevant items.
@@ 5.2 Performance comparison
Dataset Metric LightGCN LightGCN+Reg LightGCN+PC r-AdjNorm APDA PopGo Our model w/o Hetro
@@ 5.3 Ablation study
In this section, we explore the impact of different components of our proposed model on its performance. More precisely, we compare HetroFair against the situation wherein the heterophily feature weighting component is not used. We refer to this case as w/o Hetro. The results of our analysis are available in Table 2.
@@ 5.4 Hyper-parameters study
In this section, we investigate the effect of different hyper-parameters on models’ performance. Specifically, we study the impact of three hyper-parameters: the number of layers $K$, the embeddings size $d$ and $\lambda$. Due to the variety of the used datasets, we study each of the parameters over two datasets.
@@
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 4 and 5. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN and PC models drastically declines after 5 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
@@
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 6 and 7. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. Considering the NDCG metric in Figures 6 and 7, the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity.
@@
$Parameter\lambda$
Parameter $\lambda$ plays a crucial role in the accuracy-fairness trade-off. We consider the datasets Health and Beauty, and change the value of $\lambda$ from $0.05$ to $0.95$ with $0.05$ step. The results are depicted in Figures 8 and 9. Over both datasets, we observe almost the same pattern: within first few steps, the fairness metrics are in their lowest values, while the accuracy metrics are also too low. As the value of $\lambda$ increases, the performance of HetroFair increases, while PRU and RRI deteriorate until reaching to an optimal value for $\lambda$, which leads to almost stable values for all the metrics.
Figure 8: Studying the effect of $\lambda$, over the Beauty dataset.
Figure 9: Studying the effect of $\lambda$, over the Health dataset.

### v4

@@ 1 Introduction
Graph-based recommendation has become more popular in recent years due to the existence of a bipartite graph in the nature of recommender systems [5, 14, 25, 29, 30, 33]. In this bipartite graph, nodes are divided into two sets: users and items, with edges representing ratings or interactions between them. Graph neural networks (GNNs) are a powerful method for deep learning in graph-structured data, as they can capture the complex and non-Euclidean relationships between the nodes and edges of a graph [45, 10, 6, 47, 19, 22].
@@ 3 Preliminaries
In this section, we introduce preliminaries widely used in the paper, including task formulation, graph neural networks, the Light Graph Convolution model and the heterophily computation method in user-item bipartite graphs.
@@ 3.2 Graph neural networks
[…]
@@ 3.3 Light Graph Convolutional
3.4 Homophily computation in bipartite user-item graphs
To quantify homophily in the user-item bipartite graph, we use the following approach, which captures the alignment between users’ inferred preferences and the categories of the items they engage with. Each item $i\in I$ has an assigned label $\ell(i)\in\mathrm{Label}$ denoting its category. Since users lack explicit labels, we infer a dominant preference label $\ell(u)$ for each user $u\in U$ by examining the labels of the items they have interacted with. Specifically, we define:
$\ell(u)=\arg\max_{\ell\in\mathrm{Label}}\bigl{|}\{\,i\in N(u)\mid\ell(i)=\ell\,\}\bigr{|},$
where $N(u)=\{\,i\in I\mid(u,i)\in E\,\}$ is the set of items user $u$ has interacted with, and $\ell(i)$ denotes the label of item $i$. We then define a compatibility function $\mathrm{compat}(u,i)$ that indicates whether an interaction $(u,i)$ is homophilic:
$\mathrm{compat}(u,i)=\begin{cases}1,&\text{if }\ell(u)=\ell(i),\\ 0,&\text{otherwise}.\end{cases}$
Using this compatibility function, we compute the homophily score $H$ as the average compatibility across all valid edges (i.e., edges where both user and item labels are defined):
$H=\frac{1}{\lvert E^{\prime}\rvert}\sum_{(u,i)\in E^{\prime}}\mathrm{compat}(u,i),$
where $E^{\prime}\subseteq E$ contains only those interactions for which both $\ell(u)$ and $\ell(i)$ are defined. The homophily score $H$ offers insight into the structure of the interaction graph:
  • $H\approx 1$: Strong homophily—users predominantly interact with items that match their inferred preferences.
  • $H\approx 0$: Strong heterophily—users interact with items outside their dominant label.
This metric enables us to distinguish between homophilic and heterophilic datasets and to evaluate model performance accordingly.
@@ 4.2 Fairness-aware attention
$h_{u}^{\left(k\right)}=\sum_{i\in N\left(u\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}\times\delta\times\sigma(h_{u}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)})}h_{i}^{\left(k-1\right)},$ (8)
$h_{i}^{\left(k\right)}=\sum_{u\in N\left(i\right)}\frac{1}{\sqrt{d_{u}}\sqrt{d_{i}}\times\delta\times\sigma(h_{i}^{\left(k-1\right)}\cdot h_{u}^{\left(k-1\right)})}h_{u}^{\left(k-1\right)},$ (9)
where $\sigma\left(\cdot\right)$ is the sigmoid function. Parameter $\delta$ plays a pivotal role in governing the extent to which the item importance for user $u$ is either amplified or diminished. The sigmoid function transforms the dot product to a normalized range of $\left[0,1\right]$. Therefore, the result of the sigmoid function for long-tail items will be close to zero, and for short-head items, it will be close to one. By incorporating the sigmoid of dot products along with the above-mentioned normalization term, long-tail items achieve a smaller denominator than short-head items. This ensures that larger embeddings, which are more similar to each other, are divided by larger values and shrink more, while smaller embeddings are divided by smaller values and thus shrink less. Therefore messages from long-tail items (such as $i$) receive a higher importance coefficient than those from short-head items (such as $i^{\prime}$). Consequently, this augmentation enhances the similarity between users and long-tail items, compared to the previous case. Conversely, messages from short-head items are propagated with lower coefficients.
@@ 4.3 Heterophily feature weighting
$w_{ui}^{\left(k\right)}=\delta\times\sigma\left(\left(h_{u}^{\left(k-1\right)}\cdot h_{i}^{\left(k-1\right)}\right)W^{\left(k\right)}\right),$ (10)
@@
1:Input: Graph $G\left(V,E\right)$; input features matrix $X\in\mathbb{R}^{|V|\times d}$; number of layers $K$; weight matrices $W^{\left(k\right)}\in\mathbb{R}^{1\times d},\forall k\in\{1,\ldots,K\}$; parameter $\delta\cdot$
@@
8:     $w_{vi}^{\left(k\right)}\leftarrow\delta\times\sigma(\begin{bmatrix}s\end{bmatrix}\times W^{\left(k\right)})$ $\triangleright$ By $\begin{bmatrix}s\end{bmatrix}$, we mean the $1\times 1$ matrix, consisting of scalar $s$.
@@
4.4 Model training
[…]
4.5 The algorithm
Algorithm 2 One batch training process.
1:Input: Graph $G\left(V,E\right)$, the number of layers $K$.
2:Output: Updated model parameters.
[…]
4:for $k=1$ to $K$ do
6:end for
[…]
11:  Sample a negative item $j$ uniformly at random from $I\setminus O_{u}^{+}$
12:  $\tilde{y}_{ui}\leftarrow H\left[u\right]\cdot H\left[i\right]$
13:  $\tilde{y}_{uj}\leftarrow H\left[u\right]\cdot H\left[j\right]$
[…]
16:end for
17:return $\Theta$
@@ 4.5 The algorithm
Memory complexity
In addition to runtime, memory consumption is an important factor for the scalability of GNN-based recommender systems. We analyze HetroFair’s memory complexity from two perspectives: static parameter storage and dynamic training footprint. Theoretically, the memory required for our model’s parameters is comparable to that of efficient baselines such as LightGCN. The additional feature-specific weight matrices $W^{(k)}$ are shared across all nodes and scale only with the embedding dimension $d$, contributing a negligible $O(d)$ overhead to the parameter space. The primary static memory cost remains the storage of node embeddings and the graph structure, which is $O(|V|d+|E|)$, consistent with standard GNNs. However, it is important to distinguish this from the peak memory footprint during model training. Empirically, HetroFair’s training phase consumes more memory than LightGCN. This increase is not due to additional parameters but rather to the computational overhead of our fairness-aware attention mechanism. Computing dynamic, per-edge attention weights during message passing requires storing intermediate values for backpropagation, which is more memory-intensive than the sparse matrix multiplications used in LightGCN.
We argue that the additional training cost represents a reasonable trade-off given the significant and consistent gains in both fairness and recommendation accuracy that our model delivers across six datasets. Moreover, this memory overhead is confined primarily to the training phase. During inference, when user and item embeddings can be pre-computed and served, the memory requirements are substantially lower and once again comparable to those of other GNN-based recommender systems.
@@ 5.1.1 Datasets
We examine the algorithms over six well-known datasets: Epinions [^1] https://snap.stanford.edu/data/soc-Epinions1.html [44], Amazon-Movies [^2] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html [13], Amazon-Electronics2 [13], Amazon-Health2 [13], Amazon-Beauty2 [13] and Amazon-CDs2 [13]. For the amazon review datasets, we adopt a $10$-core setting, filtering out users and items with less than $10$ interactions. In the case of Epinions, we utilize the data released by APDA  [44]. Table 1 depicts the statistics of the datasets. The data highlight the diverse nature of our evaluation testbed, with significant variation in scale—from the smaller, denser Beauty and Health datasets to the larger, sparser Movies and Electronics datasets. This diversity is crucial for validation, as their density suggests that popularity bias and long-tail issues may manifest differently across them. Furthermore, the homophily rate spans a wide spectrum, from highly heterophilic (e.g., Movies at 0.14) to strongly homophilic (e.g., Health at 0.89), providing a robust environment for evaluating the effectiveness of our heterophily-aware components.
@@
Dataset #users #items #interactions Density Homophily rate
Electronics 20242 11589 347393 0.148% 0.15
Epinions 11496 11656 327942 0.245% N/A [^3] The Epinions dataset lacks the category information necessary for computing the homophily rate.
Beauty 1340 733 28798 2.93% 0.79
Health 2184 1260 55076 2% 0.89
Movies 33326 21901 958986 0.131% 0.14
CDs 15592 16184 445412 0.176% 0.34
@@ 5.1.3 Evaluation metrics
Here, $SRC\left(\cdot,\cdot\right)$ calculates Spearman’s rank correlation, $\tilde{O}_{u}^{+}$ represents items in the test profile of user $u$, $rank_{u}\left(x\right)$ signifies the ranking of item $x$ that the model predicts for user $u$, $U_{i}$ is a set of users that item $i$ already exists in their test profiles and $d_{x}$ is the degree of node $x$. While PRU investigates items’ position bias for individual users’ recommendation lists, PRI considers item-side fairness across all users. A large value of the later may result in long-tail items having limited opportunities to become popular [46]. In general, the methods’ performance in promoting fairness improves as the values of these two criteria get closer to 0.
@@
Here, $REL_{N}$ is a sorted list of the $N$ most relevant items in ascending order, $r(i)$ is the relevance of the $i^{th}$ ranked item, which is either 0 or $1$ respectively for irrelevant and relevant items. In our experiments, we set $N=20$, and this value is applied consistently across all metrics.
@@ 5.2.1 Overall performance across metrics
An important observation from Table 2 is the consistent trade-off between accuracy and fairness among the baseline models. For instance, Reg often improves fairness metrics over the standard LightGCN, but this typically comes at the cost of accuracy (lower NDCG). This behavior reflects the inherent difficulty of mitigating popularity bias without degrading recommendation relevance. In contrast, our proposed HetroFair model consistently achieves a more favorable balance across datasets: it improves fairness metrics such as PRU and PRI (bringing them closer to zero) while simultaneously enhancing accuracy metrics. This suggests that our model not only addresses item-side unfairness effectively, but also leverages the underlying graph structure to maintain or improve overall recommendation quality.
@@
Dataset Metric LightGCN Reg PC r-AdjNorm APDA PopGo Our model w/o Hetro
@@
5.2.2 Analysis by dataset homophily
A deeper analysis of our results, when viewed through the lens of the datasets’ homophily rates, reveals a compelling narrative about HetroFair’s specific strengths. Our proposed model demonstrates its most significant performance gains on datasets characterized by low to moderate homophily, where user preferences are more diverse. For instance, on highly heterophilic datasets such as Movies and Electronics, and moderately heterophilic ones like Epinions and CDs, HetroFair consistently and substantially outperforms all baseline methods across nearly every accuracy and fairness metric. This suggests that its fairness-aware attention and heterophily-centric design are particularly effective at capturing signals from diverse user-item interactions.
Conversely, on the two most homophilic datasets, Health and Beauty, the performance gap narrows as the interactions become more uniform. Even in these settings, however, HetroFair remains highly competitive. For example, on the Health dataset, our model achieves the highest accuracy scores across all three performance metrics. On the Beauty dataset, it secures the second-best performance for accuracy while delivering the top performance for the PRI fairness metric. While other models like PopGo also show strong results on certain metrics in these environments, our model’s ability to consistently deliver a leading or competitive performance highlights its robustness. This overall pattern validates our model’s core design principle: HetroFair is explicitly built to be heterophily-aware, and its capacity to excel in those conditions while remaining a top contender elsewhere confirms its effective and well-rounded architectural design.
@@ 5.3 Ablation study
In this section, we explore the impact of different components of our proposed model on its performance. More precisely, we compare HetroFair against the situation wherein the heterophily feature weighting component is not used. We refer to this case as w/o Hetro. In addition to that we test different initialization strategies for the feature-specific weights. The results of our analysis are available in Table 2 and Figure 4..
5.3.1 Heterophily feature weighting
To analyze the effect of this module, we disable the feature-specific weighting mechanism while keeping all other components unchanged, and evaluate its contribution to overall performance across accuracy and fairness metrics:
@@ 5.3.1 Heterophily feature weighting
5.3.2 Feature weight initialization
We conduct an ablation study to evaluate the robustness of the model with respect to the initialization of the trainable feature-specific weights $\mathbf{W}^{(k)}$. Specifically, we compare three initialization strategies: zero initialization, random normal initialization, and Xavier initialization (used in the main model). As illustrated in Figure 4, we assess model performance across three datasets (Epinions, Health, and Beauty) and three key metrics: NDCG, PRU, and PRI. While all three schemes yield comparable results, Xavier initialization consistently provides a more favorable balance between accuracy and fairness. Zero initialization tends to yield weaker fairness outcomes, likely due to the absence of early asymmetry in gradient flow, which limits the model’s ability to differentiate between features during early training. Random normal initialization introduces stochastic variability, but occasionally leads to less stable fairness metrics. Xavier initialization, by preserving the variance of input and output signals, avoids saturation and ensures more effective gradient propagation. These observations indicate that the choice of initialization, although not dramatically altering overall trends, can affect stability and fairness. Xavier initialization offers a stable and generalizable solution, making it a reliable choice for initializing feature weights in fairness-aware GNN models.
(a) NDCG
(b) PRU
(c) PRI
Figure 4: Performance of the model under different initialization strategies for feature-specific weights. .
@@ 5.4 Hyper-parameters study
In this section, we investigate the effect of different hyper-parameters on models’ performance. Specifically, we study the impact of three hyper-parameters: the number of layers $K$, the embeddings size $d$ and $\delta$. Due to the variety of the used datasets, we study each of the parameters over two datasets.
@@
To study the impact of the number of GNN layers, we select two datasets Beauty and Epinions, and investigate the potential effect of over-smoothing over them. We evaluate the impact of the number of layers by ranging them from 1 to 6, and use NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 5 and 6. For the Epinions dataset, our model outperforms all the baselines across all the metrics. In all the methods, NDCG increases until reaching the optimal point and then starts to decline due to the over-smoothing phenomenon. As the number of layers increases, PRU and PRI metrics decrease, indicating that increasing the number of layers up to a certain point can alleviate the unfairness caused by popularity bias. The effect of over-smoothing becomes less noticeable due to the use of the weighted sum approach for the final node representation. On the other hand in the Beauty dataset, although we use the same weighted sum approach, the effect of the over-smoothing problem is evident from the initial stages of the models. The performance of the LightGCN and PC models drastically declines after 5 GNN layers. The performance of HetroFair decreases too, but with a smaller slope and it still outperforms all the baselines over the Beauty dataset. This shows its robust performance across different settings.
APDAHetro-FairLightGCNPCPopGoRegr-AdjNorm
(a) NDCG
(b) PRU
(c) PRI
@@
The embeddings dimension is an important factor in graph-based recommendation systems, representing the latent factors size of users and items. We consider two datasets CDs and Epinions, and compare the performance of the models across three embeddings sizes: $64$ (similar to LightGCN and r-AdjNorm), $128$ (the default value for HetroFair) and $256$ (the used value by APDA). We utilize NDCG as the accuracy metric and PRU and PRI as the fairness metrics. The results are presented in Figures 7 and 8. As can be seen in the figures, over both datasets and by increasing the embeddings dimension, NDCG of HetroFair increases, too. In other words, in our model NDCG has positive correlation with the embeddings dimension. For all values of the embeddings dimension, HetroFair outperforms the other methods, in terms of both accuracy and fairness measures. It can also be seen in the figures that the dependence of PRI on the dimension size is negligible. Considering the NDCG metric in Figures 7 and 8, the plots exhibit steeper slopes when the dimensions change from $64$ to $128$, compared to the cases where the dimensions change from $128$ to $264$. This means that the rates of accuracy changes decrease from 128 dimensions and beyond. According to our time complexity analysis in Section 4, time complexity of HetroFair increases linearly with the dimension of the embeddings. Therefore, considering $128$ as the default dimension for embeddings seems to be a good trade-off between accuracy and complexity.
@@
$Parameter\delta$
Parameter $\delta$ plays a crucial role in the accuracy-fairness trade-off. We consider the datasets Health and Beauty, and change the value of $\delta$ from $0.05$ to $0.95$ with $0.05$ step. The results are depicted in Figures 9 and 10. Over both datasets, we observe almost the same pattern: within first few steps, the fairness metrics are in their lowest values, while the accuracy metrics are also too low. As the value of $\delta$ increases, the performance of HetroFair increases, while PRU and RRI deteriorate until reaching to an optimal value for $\delta$, which leads to almost stable values for all the metrics.
Figure 9: Studying the effect of $\delta$, over the Beauty dataset.
Figure 10: Studying the effect of $\delta$, over the Health dataset.